OpenStudy (anonymous):

How much will $1000 deposited in an account earning 7% interest compounded annually be worth in 20 years?

OpenStudy (solomonzelman):

This is the formula you need: \(\Large\color{blue}{ \displaystyle {\rm A}= {\rm P}\left(1+\frac{ {\rm r} }{ {\rm n} }\right)^{{\rm n} \times {\rm t}} }\) \(\Large\color{blue}{ \displaystyle {\rm r} }\) - the interest rate (written as a decimal) \(\Large\color{blue}{ \displaystyle {\rm n} }\) - the number of times that interest is compounded per year. \(\Large\color{blue}{ \displaystyle {\rm t} }\) - number of years (of investment). \(\Large\color{blue}{ \displaystyle {\rm P} }\) - the initial amount (the original deposite that you made into the bank). \(\Large\color{blue}{ \displaystyle {\rm A} }\) - the amount that you will get after this investment.

OpenStudy (solomonzelman):

P = 1000 (that is what you begin with) r = 0.07 (because your interestrate is 7%, and that is 0.07 as a decimal) n = 1 (because it says "compunded annually") t = 20 (because it is asking you for the total amount after 20 years of investment)

OpenStudy (solomonzelman):

All you have to do is to plug your numbers and calculate the value of the investment (the value of A). If you have questions, ask.

OpenStudy (anonymous):

Ok thanks

OpenStudy (solomonzelman):

yw